. 25% of all Ethiopian households had VCR in 2002. If 500 households are randomly selected from the population, what is the probability that 100 or more households have VCR?



1
Expert's answer
2021-12-24T08:56:09-0500

If "X" is a binomial random variable with mean "\u03bc = np" and variance "\u03c3^2 = npq," then the limiting form of the distribution of "Z=\dfrac{X-np}{\sqrt{npq}}," as "n\to \infin," is the standard normal distribution "n(z;0,1)."

In practice, the approximation is adequate provided that both "np\geq 10" and

"nq\geq 10."

Given "n=500, p=0.25, q=1-p=0.75"


"np=500(0.25)=125\geq 10"

"nq=500(0.75)=375\geq 10"

The approximation can safely be applied.

Use Normal Distribution with Continuity Correction


"P(X\geq100)\approx P(X>99.5)=1-P(X\leq99.5)"

"=1-P(Z\leq\dfrac{99.5-125}{\sqrt{500(0.25)(0.75)}})"

"\approx1-P(Z\leq-2.63363)\approx0.995776"